An undergraduate business student has purchased a laptop computer for use during exams. This laptop is perfectly reliable except for two parts: its microchip, which has a MTBF = 10 hours and its battery, which has a MTBF = 20 hours. Assuming that a new battery has just been installed, what is the probability that the LAPTOP will FAIL during a 2 hour exam?
Answer #1

Answer: 0.98

Explanation: failure probability to fail in 2 hour due to chip = 2/10 = 1/5
failure probability to fail in 2 hour due to battery = 2/20 = 1/10
combined failure probability to fail in 2 hour = 1/5* 1/10
now reliable performance probability = (1 - (1/5* 1/10)) = 0.98

Please ask, if you have any doubts through the comment section. Do rate the answer